# Robust Superradiance and Spontaneous Spin Ordering in Disordered Waveguide QED

### v1

@@ Appendix D Gaussian-Distributed Positions
Figure 13: Superradiance with Gaussian-distributed positions. Scaling of (a) superradiant rate $\mathcal{R}/N^{2}$ and (b) burst time $t_{\star}N/\log(N)$, for Gaussian-distributed positions with disorder strengths $\Theta=\pi/2,\pi,3\pi/2$, showing convergence to $N^{2}$ scaling with different finite-size scalings. We find that the data can be fitted with the function $N^{2}(r_{0}+r_{1}N^{-1})$ for $\Theta=\pi/2$ (dashed line), and $N^{2}(r_{0}+r_{1}^{\prime}N^{-1/2})$ for $\Theta=\pi$ and $\Theta=3\pi/2$ (solid lines). Results are averaged over $10^{3}$ trajectories.
@@
For completeness, we here consider the case of Gaussian position fluctuations instead of the uniform distribution considered in the main text. Specifically, let the propagation phases $\{\xi_{j}\}$ obey a Gaussian distribution $\mathcal{N}(0,\Theta^{2}/4)$, with standard deviation $\Theta/2$ for comparison with the uniform case considered above. Then we know $\Re e^{-i2\xi}$ and $\Im e^{-i2\xi}$ have the following means and variances respectively:
$\tilde{\mu}_{a}$ $=e^{-\Theta^{2}/2},$ $\tilde{\sigma}_{a}^{2}$ $=\frac{1}{2}(1-2e^{-\Theta^{2}}+e^{-2\Theta^{2}}),$
$\tilde{\mu}_{b}$ $=0,$ $\tilde{\sigma}_{b}^{2}$ $=\frac{1}{2}(1-e^{-2\Theta^{2}}).$ (54)
@@
Then, in the large $N$ limit, $c$ obeys a Gaussian distribution with $\Re c\sim N(\tilde{\mu}_{a},\tilde{\sigma}_{a}^{2}/N)$, $\Im c\sim N(\tilde{\mu}_{b},\tilde{\sigma}_{b}^{2}/N)$. Following the same procedure as in Appendix A, we know, in the limit $N\rightarrow\infty$,
$\Gamma_{\pm}\rightarrow\frac{1}{2}\gamma N(1\pm\tilde{\mu}_{a})=\frac{1}{2}\gamma N(1\pm e^{-\Theta^{2}/2}),$ (55)
which still gives upper bound $\mathbb{R}_{>}\sim N^{2}$ for all disorder strength.
@@
For small $\Theta$, $\tilde{\mu}_{a}\sim 1$, the expansion in Eq. (47) gives
$E[|c|]=\tilde{\mu}_{a}+\frac{1}{2\tilde{\mu}_{a}}E[\delta_{b}^{2}]\sim\tilde{\mu}_{a}+O(N^{-1}).$ (56)
That is,
$E[\Gamma_{\max}]=\frac{1}{2}\gamma N\big(1+\tilde{\mu}_{a}+O(N^{-1})\big).$ (57)
@@
The finite-size correction is of order $1/N$ for $\Gamma_{\max}/N$. While for large $\Theta$ and finite $N$, we can have $\tilde{\mu}_{a}\ll\delta_{a}$ i.e. $\Theta\gg\sqrt{\log(N)}$ in Eq. (46). That is,
$E[|c|]\approx E[\sqrt{(\delta_{a}^{2}+\delta_{b}^{2})}]\sim E[|\delta_{a/b}|]\sim N^{-1/2},$ (58)
which gives slow finite-size scaling
$E[\Gamma_{\max}]\approx\frac{1}{2}\gamma N\big(1+O(N^{-1/2})\big).$ (59)
@@
In Fig. 13, we show $\mathcal{R}_{\star}/N^{2}$ vs. NN for a Gaussian distribution with $\Theta=\pi/2,\pi,3\pi/2$. For $\Theta=\pi/2$, $E[|c|]\approx 0.29$, Eq. (57) predicts a finite size scaling $\sim O(N^{-1})$, which agrees with our numerical data. As predicted in Eq. (59), $\Theta=\pi$ and $3\pi/2$ gives $E[|c|]\approx 7\times 10^{-3}$ and $1.1\times 10^{-5}$ respectively, which leads to a finite size scaling $\sim O(N^{-1/2})$ as observed in numerical simulations with $N\leq 400$.
Following Sec. V.1, we can derive the upper bound $\mathbb{R}_{<}$ for Gaussian-distributed positions. We see
$\mathbb{R}_{<}=\frac{\gamma N^{2}}{8}(1+e^{-\Theta^{2}}),$ (60)
@@
which predicts an $N^{2}$ scaling as confirmed in Fig. 13. Similar to the case with homogeneous distribution, $\mathbb{R}_{<}$ always has finite-size correction $\sim N^{-1}$, but does not capture the transient $N^{-1/2}$ scaling.
Appendix E Neglecting the Hamiltonian
In Fig. 14, we show results of DTWA simulations without the Hamiltonian $H$. Since the jump operators are the same as those of the original problem, the product state ansatz thus predicts the same upper bound $\mathbb{R}$. Indeed, it can be seen that removing $H$ only increases the prefactor of the $N^{2}$ scaling but not the scaling itself, confirming the prediction in Sec. V.1.
Figure 14: Comparison of superradiant scaling with (solid lines) and without (dashed lines) the coherent Hamiltonian $H$. (a) Superradiant rate $\mathcal{R}_{\star}/N^{2}$ and (b) burst time $t_{\star}$ as a function of $N$, under different disorder strengths. Removing $H$ changes only the prefactor of the $N^{2}$ scaling, not the scaling itself.
Appendix F A Loose Upper Bound
@@
Since $\hat{\mathcal{R}}\geq 0$, we know $\operatorname{Tr}(\hat{\mathcal{R}})=\operatorname{Tr}(\hat{\mathcal{R}}_{1})+\operatorname{Tr}(\hat{\mathcal{R}}_{2})\geq 0$. From $\operatorname{Tr}(\hat{\mathcal{R}}_{1})=\gamma N$, we know $\operatorname{Tr}(\hat{\mathcal{R}}_{2})<\gamma N$, which bounds maximal eigenvalue of $\hat{\mathcal{R}}_{2}$ by $\gamma N$ from above. That is, we can further have [64]
@@
$\begin{split}\lambda_{\max}(\hat{\mathcal{R}})\leq\frac{3}{2}\Gamma_{\max}N-\frac{1}{2}\gamma N\rightarrow\frac{3\gamma}{4}N^{2}(1+\frac{\sin\Theta}{\Theta}),\end{split}$ (70)
@@
Appendix G Time Evolution of Left- and Right-Emission Rates
@@
In complement to Fig. 9, Fig. 15(a) shows the time evolution of the left-right rate difference $(\mathcal{R}_{\text{R}}-\mathcal{R}_{\text{L}})$ in QSDMF trajectories for $\Theta=2\pi$. It can be seen that the rate difference deviates from 0 and diffuses as time increases, resulting in symmetrical fluctuations around the mirror-symmetric emission. As shown in Fig. 15(b), this fluctuation peaks at $t_{\star}$ and increases with disorder strength. Note that the trajectories can change their preferred emission direction during the decay, which means that the emergence of the two spin orderings is not a bifurcation-like process.
Figure 16: Spontaneous spin ordering in regular lattices. (a) and (b) $p(\Delta\xi,\Delta\phi)$ of the relative propagating phase $\Delta\xi$ and relative spin orientation $\Delta\phi$ for $k_{0}d=\pi/3$ and $k_{0}d=\pi/2$, respectively. Note that each slice is infinitely narrow, and the plotting shows a finite width for data presentation. (c) and (d) are the corresponding distribution of $\Delta\phi$ at a fixed distance $\Delta z$. See the Supplemental Material [65] for results on other lattice constants. Simulations use $N=400$ atoms and $10^{3}$ trajectories.
Appendix H Spontaneous Spin Ordering in a Regular Lattice
Here, besides the disordered case considered in Sec. V.2, we would like to show that the spontaneous spin ordering similarly exists for a regular lattice with a lattice constant $d$. The upper bound can be written as
For $k_{0}d\leq\pi/(2N)$, we know $\left(i-j\right)k_{0}d\leq\pi/2$, which makes $\{\phi_{i}=\phi_{j}\}$ the optimal configuration. For a larger $k_{0}d$, we expect the approximate bound $\mathbb{R}_{<}$ to work well as in Sec.V.2, which predicts spins ordering according to $\{\phi_{j}=\phi_{1}+(j-1)k_{0}d\}$ or $\{\phi_{j}=\phi_{1}-(j-1)k_{0}d\}$. Then a lower estimate of the upper bound is given by
Specifically, we expect ordering patterns similar to spin waves, with a period of $2\pi/(k_{0}d)$. In Fig. 16, we show the distribution of the relative angles $\Delta\phi$ at distance $\Delta z$ between a pair of spins, for different lattice constants $d$. There are peaks around $\Delta\phi\approx\pm\Delta zk_{0}d$ as predicted from the product state ansatz. It can be seen that the ferromagnetic and antiferromagnetic orderings are more pronounced than other configurations. Interestingly, distributions $\Delta\phi\approx\pm\pi/2$ are almost flat compared with other peaks. This is because, in Eq. (71), the terms with $(i-j)k_{0}d\approx\pm\pi/2$ vanish, which makes the value of $\Delta\phi$ arbitrary in the optimal configuration.
@@ Appendix H Spontaneous Spin Ordering in a Regular Lattice
I Irrelevance of Detuning in DTWA
In the DTWA formalism, after the unitary transformation for the rotating frame, the EOMs (53) in Appendix B become
@@ I Irrelevance of Detuning in DTWA
with $\tilde{s}_{j}^{\pm}(t)=s_{j}^{\pm}(t)e^{\mp i\delta\omega_{j}t}$. Notice that the detuning gives corrections $\sim\Delta_{\omega}t\sim(\Delta_{\omega}/\gamma)\ln(N)/N$ to the EOMs, which become negligible for $N/\ln(N)\gg\Delta_{\omega}/\gamma$. Indeed, in Fig. 13, we can see that detuning does not change the superradiant rate when $N$ is sufficiently large. Since DTWA is accurate for this problem, this also indicates that the required cumulant order $C\ll N/\ln(N)$ for superradiance. This is numerically confirmed in Appendix C that $C=2$ or $3$ is sufficient to give accurate results in this setting.
J Supplementary Plots
Figure S1: Changing of spin ordering as disorder strength $\Theta$ increases, in complement to Fig. 8 and 9. Upper: Joint distribution $p(\Delta\xi,\Delta\phi)$ at superradiant time $t_{\star}$ for 1000 trajectories with $N=400$. Lower: Joint distribution of $(\mathcal{R}_{\text{R}},\mathcal{R}_{\text{L}})$ for 1000 trajectories with $N=400$. The clockwise and anticlockwise spin orderings gradually become more and more evident as $\Theta$ increases. Then, we can see that the fluctuation around the mirror-symmetric emission also increases gradually as the disorder strength increases.
Figure S3: (a) Probability distribution of relative propagating phases $\Delta\xi=(\xi_{i}-\xi_{j})\bmod 2\pi$, which is a wrapped triangular distribution. (b) Probability distribution of relative spin angles $\Delta\phi=(\phi_{i}-\phi_{j})\bmod 2\pi$. When $\Theta=0$, the system is permutation-invariant in QSDMF trajectories, and all spins align along the same direction, resulting in $p(\Delta\phi)=\delta(0)$ as expected. As $\Theta$ increases, the relative propagating phase $\Delta\xi$ becomes more uniformly distributed in $(-\pi,\pi)$. Accordingly, $p(\Delta\phi)$ also becomes more evenly distributed, meaning the global spin alignment of the atoms weakens. Instead, a new spin ordering emerges, with spins aligning according to their position-dependent phases as revealed by the joint probability distribution $p(\Delta\xi,\Delta\phi)$ shown in Sec. V.2. Additionally, we can observe that, under strong disorder, there is a weak tendency for alignment and anti-alignment in the distribution $p(\Delta\phi)$. This can be attributed to the fact that the joint distribution $p(\Delta\xi,\Delta\phi)$, shown in Fig. 8, has less spreading around $\Delta\xi\approx 0$ and $\pi$. For similar reasons as in Appendix H, the ordering around $\Delta\phi\approx\Delta\xi\approx\pm\pi/2$ is weaker than ordering in other directions.
Figure S4: In complement to Fig. 11, (a) Error of DTWA for $N=5,10,15$ atoms coupled with one lossy cavity, (b) Burst time $t_{\star}$ vs. NN for atoms coupled with two cavities, for $\kappa/(\gamma N)=1,2,5,10,50$ (from bottom to top), and (c) The roughly linear relation between $\gamma t_{\star}N\ln(N)$ and $\gamma N/\kappa$ for $N=400$. $t_{\star}\sim\ln(N)/(\gamma N)$ scaling persists in the weakly non-Markovian regime, with the prefactor decreasing to 11 as $\kappa$ increases.

### v2

@@ Appendix D Gaussian-Distributed Positions
Figure 13: Superradiance with Gaussian-distributed positions. Scaling of (a) the superradiant rate $\mathcal{R}/(\gamma N^{2})$ and (b) the burst time $t_{\star}\gamma N/\log(N)$, for Gaussian-distributed positions with disorder strengths $\Theta=\pi/2,\pi,3\pi/2$. The plots demonstrate a convergence to an $N^{2}$-scaling, but with different finite-size corrections. We find that the data can be fitted with the function $N^{2}(r_{0}+r_{1}N^{-1})$ for $\Theta=\pi/2$ (dashed line), and $N^{2}(r_{0}+r_{1}^{\prime}N^{-1/2})$ for $\Theta=\pi$ and $\Theta=3\pi/2$ (solid lines). Results are averaged over $\mathcal{N}=10^{3}$ trajectories.
@@
For completeness, we here consider the case of Gaussian position fluctuations instead of the uniform distribution considered in the main text. Specifically, let the propagation phases $\{\xi_{j}\}$ obey a Gaussian distribution $\mathcal{N}(0,\Theta^{2}/4)$, with standard deviation $\Theta/2$ for comparison with the uniform case considered above. Then, the quantities $\Re e^{-i2\xi}$ and $\Im e^{-i2\xi}$ have the following mean values and variances:
$\tilde{\mu}_{a}$ $=e^{-\Theta^{2}/2},$ $\tilde{\sigma}_{a}^{2}$ $=\frac{1}{2}\left(1-2e^{-\Theta^{2}}+e^{-2\Theta^{2}}\right),$
$\tilde{\mu}_{b}$ $=0,$ $\tilde{\sigma}_{b}^{2}$ $=\frac{1}{2}\left(1-e^{-2\Theta^{2}}\right).$ (55)
@@
Consequently, in the large $N$ limit, $c$ obeys a Gaussian distribution with $\Re c\sim N(\tilde{\mu}_{a},\tilde{\sigma}_{a}^{2}/N)$ and $\Im c\sim N(\tilde{\mu}_{b},\tilde{\sigma}_{b}^{2}/N)$. Following the same procedure as in Appendix A, we obtain for $N\rightarrow\infty$,
$\Gamma_{\pm}\rightarrow\frac{1}{2}\gamma N(1\pm\tilde{\mu}_{a})=\frac{1}{2}\gamma N\left(1\pm e^{-\Theta^{2}/2}\right).$ (56)
@@
For small $\Theta$, we have $\tilde{\mu}_{a}\sim 1$, and using the expansion in Eq. (47), we obtain the scaling
$E[\Gamma_{+}]=\frac{1}{2}\gamma N\big(1+\tilde{\mu}_{a}+O(N^{-1})\big),$ (57)
@@
with a finite-size correction of order $1/N$ for $\Gamma_{\max}/N$. For large $\Theta$ and finite $N$, i.e., $\Theta\gg\sqrt{\log(N)}$, we encounter the relation $0<\tilde{\mu}_{a}\ll\delta_{a}$, and, according to Eq. (46),
$E[|c|]\approx E[\sqrt{(\delta_{a}^{2}+\delta_{b}^{2})}]\sim E[|\delta_{a/b}|]\sim N^{-1/2}.$ (58)
Therefore, under this condition we find a slow finite-size scaling
$E[\Gamma_{+}]\approx\frac{1}{2}\gamma N\left(1+O(N^{-1/2})\right).$ (59)
@@
In Fig. 13, we plot $\mathcal{R}_{\star}/(\gamma N^{2})$ as a function of $N$ for Gaussian distributions with $\Theta=\pi/2,\pi,3\pi/2$. For $\Theta=\pi/2$, $E[|c|]\approx 0.29$, Eq. (57) predicts a finite-size scaling $\sim O(N^{-1})$, which agrees with our numerical data. As predicted in Eq. (59), $\Theta=\pi$ and $\Theta=3\pi/2$ give $E[|c|]\approx 7\times 10^{-3}$ and $E[|c|]\approx 1.1\times 10^{-5}$, respectively, which leads to a finite-size scaling $\sim O(N^{-1/2})$ as observed in numerical simulations with $N\leq 400$.
Following Sec. V.1, we can derive the upper bound
$\mathbb{R}_{<}=\frac{\gamma N^{2}}{8}\left(1+e^{-\Theta^{2}}\right),$ (60)
@@
for a Gaussian distribution, which tightly bounds the numerical data from above. Similar to the case with homogeneous distribution, $\mathbb{R}_{<}$ always has finite-size correction $\sim N^{-1}$, but does not capture the transient $N^{-1/2}$ scaling.
Appendix E A Loose Upper Bound
@@
Since $\hat{\mathcal{R}}\geq 0$, we know $\braket{\hat{\mathcal{R}}}=\braket{\hat{\mathcal{R}}_{1}}+\braket{\hat{\mathcal{R}}_{2}}\geq 0$. From $\operatorname{Tr}(\hat{\mathcal{R}}_{1})=\gamma N$, we know $\braket{\hat{\mathcal{R}}_{2}}\leq\gamma N$, which bounds maximal eigenvalue of $\hat{\mathcal{R}}_{2}$ by $\gamma N$ from above. That is, we can further have [68]
@@
$\begin{split}\lambda_{\max}(\hat{\mathcal{R}})\leq\frac{3}{2}\Gamma_{\max}N-\frac{1}{2}\gamma N\rightarrow\frac{3\gamma}{4}N^{2}\left(1+\frac{\sin\Theta}{\Theta}\right),\end{split}$ (70)
@@
Appendix F Neglecting the Hamiltonian
In Fig. 14, we show results of DTWA simulations without the Hamiltonian $H$. Since the jump operators are the same as those of the original problem, the product state ansatz thus predicts the same upper bound $\mathbb{R}$. Indeed, it can be seen that removing $H$ only increases the prefactor of the $N^{2}$-scaling but not the scaling itself, confirming the prediction of Sec. V.1. Since the dissipator only depends on $\{z_{j}\bmod{2\pi}\}$, the conclusion in this study therefore holds for any shift of locations by integer multiples of wavelength $\lambda_{0}$.
Figure 14: Comparison of superradiant scaling with (solid lines) and without (dashed lines) the coherent Hamiltonian $H$. (a) Plot of the superradiant rate $\mathcal{R}_{\star}/(\gamma N^{2})$ and (b) the burst time $t_{\star}$ as a function of $N$ and for different disorder strengths. The Hamiltonian part of the evolution only slightly changes the prefactor of the $N^{2}$-scaling, but not the scaling itself.
Appendix G Spontaneous Spin Ordering in a Regular Lattice
Figure 15: Spontaneous spin ordering in regular lattices. (a) and (b) Plot of the distribution $p(\Delta\xi,\Delta\phi)$ of the relative propagating phase $\Delta\xi$ and relative spin orientation $\Delta\phi$ for $k_{0}d=\pi/3$ and $k_{0}d=\pi/2$, respectively. Note that for a regular array, each slice in $\Delta z$ is infinitely narrow, but has been broadened for better visibility. The plots in (c) and (d) show the corresponding probability distribution function (PDF) of $\Delta\phi$ at a fixed distance $\Delta z$. The color bar for (a) and (b) shows the corresponding PDF in (c) and (d). See the Supplemental Material [60] for similar results for other lattice constants. Simulations use $N=400$ atoms and $10^{3}$ trajectories.
In addition to the disordered atomic arrays considered in Sec. V.2, in this appendix we show that a similar spontaneous spin ordering exists for regular lattices with a lattice constant $d$. In this case, the upper bound can be written as
For $k_{0}d\leq\pi/(2N)$, we know $\left(i-j\right)k_{0}d\leq\pi/2$, which makes $\{\phi_{i}=\phi_{j}\}$ the optimal configuration. As discussed in in Sec.V.2, for a larger value of $k_{0}d$, we expect the approximate bound $\mathbb{R}_{<}$ to provide an accurate estimate for the maximal decay rate. It predicts an ordering of the atomic dipoles according to $\{\phi_{j}=\phi_{1}+(j-1)k_{0}d\}$ or $\{\phi_{j}=\phi_{1}-(j-1)k_{0}d\}$. Then a lower estimate of the upper bound is given by
Specifically, we expect ordering patterns similar to spin waves, with a period of $2\pi/(k_{0}d)$. In Fig. 15, we show the distribution of the relative angles $\Delta\phi$ at distance $\Delta z$ between a pair of spins, for different lattice constants $d$. There are peaks around $\Delta\phi\approx\pm\Delta zk_{0}d$ as predicted from the product state ansatz. It can be seen that the ferromagnetic and antiferromagnetic orderings are much more pronounced than other configurations. Interestingly, distributions $\Delta\phi\approx\pm\pi/2$ are almost flat compared to other cases. This is because, in Eq. (71), the terms with $(i-j)k_{0}d\approx\pm\pi/2$ vanish, which makes the value of $\Delta\phi$ arbitrary in the optimal configuration.
Appendix H Time Evolution of Left- and Right-Emission Rates
@@
As a complement to Fig. 9, Fig. 16(a) shows the time evolution of the left-right emission asymmetry $(\mathcal{R}_{\text{R}}-\mathcal{R}_{\text{L}})$ for individual QSDMF trajectories for $\Theta=2\pi$. It can be seen that most trajectories exhibit an asymmetry in the emission rate with a preference for either one or the other side. However, this asymmetry is not completely fixed during the decay and some of the trajectories change their preferred direction of emission during the decay process. As shown in Fig. 16(b), relative strength of the asymmetry peaks at $t_{\star}$ and increases with disorder strength.
Appendix I Boundary Effects
Figure 17: Probability distribution of the decay-rate asymmetry obtained from QSDMF trajectories, after removing the data from the last $n_{\text{B}}=0,10,20,30,40,50$ atoms at each boundary, i.e. keeping only the contribution from the bulk $(N-2n_{B})$ atoms. As $n_{\text{B}}$ increases, only contributions from the bulk remain. The two peaks correspond to the two spontaneous spin-ordering configurations, while the boundary atoms introduce only minor modifications to the distribution. Simulations for this plot have been performed with $N=400$, $\Theta=2\pi$, and using $5\times 10^{4}$ trajectories.
In a recent work [69], the authors investigate closely related spin-ordering effects in the steady states of actively pumped atomic arrays and observe a strong influence of the boundaries: spins at the left (right) boundary tend to favor the anticlockwise (clockwise) configurations. As a result, boundary atoms contribute approximately symmetrically to emission into the left- and right-propagating modes.
To quantify the influence of boundary atoms on spin ordering in the current setting, Fig. 17 shows the probability distribution of the decay-rate asymmetry $\mathcal{A}=(\mathcal{R}_{\text{R}}-\mathcal{R}_{\text{L}})/(\mathcal{R}_{\text{R}}+\mathcal{R}_{\text{L}})$, which is computed after artificially removing the contributions of $n_{\text{B}}$ atoms from each boundary. The distribution exhibits two pronounced peaks, corresponding to the two bulk spin-ordering configurations $\rho^{\pm}$. As $n_{\text{B}}$ increases, the contribution of the boundary atoms is progressively suppressed, leading to a slight enhancement of the two peaks, but no significant changes. These results show that while boundary atoms do exhibit behavior distinct from that of the bulk, their effect on the decay-rate asymmetry is minor. The dominant contribution arises from the bulk atoms, which realize the two spontaneous spin-ordering configurations predicted in Sec. V.2.
@@ Appendix I Boundary Effects
J Supplementary Plots
  • Fig. S4 shows probability distributions of the spin angle $\theta$ at superradiant times, for different disorder strength $\Theta$.
Figure S1: Changing of spin ordering as disorder strength $\Theta$ increases, in complement to Fig. 8 and 9. Upper: Joint distribution $p(\Delta\xi,\Delta\phi)$ at superradiant time $t_{\star}$ for $10^{3}$ trajectories with $N=400$. Lower: Joint distribution of $(\mathcal{R}_{\text{R}},\mathcal{R}_{\text{L}})$ for $10^{3}$ trajectories with $N=400$. The clockwise and anticlockwise spin orderings gradually become more and more evident as $\Theta$ increases. Then, we can see that the fluctuation around the mirror-symmetric emission also increases gradually as the disorder strength increases.
Figure S3: (a) Probability distribution of relative propagating phases $\Delta\xi=(\xi_{i}-\xi_{j})\bmod 2\pi$, which is a wrapped triangular distribution. (b) Probability distribution of relative spin angles $\Delta\phi=(\phi_{i}-\phi_{j})\bmod 2\pi$, for $N=400$ and $10^{3}$ trajectories. When $\Theta=0$, the system is permutation-invariant in QSDMF trajectories, and all spins align along the same direction, resulting in $p(\Delta\phi)=\delta(0)$ as expected. As $\Theta$ increases, the relative propagating phase $\Delta\xi$ becomes more uniformly distributed in $(-\pi,\pi)$. Accordingly, $p(\Delta\phi)$ also becomes more evenly distributed, meaning the global spin alignment of the atoms weakens. Instead, a new spin ordering emerges, with spins aligning according to their position-dependent phases as revealed by the joint probability distribution $p(\Delta\xi,\Delta\phi)$ shown in Sec. V.2. Additionally, we can observe that, under strong disorder, there is a weak tendency for alignment and anti-alignment in the distribution $p(\Delta\phi)$. This can be attributed to the fact that the joint distribution $p(\Delta\xi,\Delta\phi)$, shown in Fig. 8, has less spreading around $\Delta\xi\approx 0$ and $\pi$. For similar reasons as in Appendix G, the ordering around $\Delta\phi\approx\Delta\xi\approx\pm\pi/2$ is weaker than ordering in other directions.
Figure S4: Probability distribution of each spin angle $\theta_{i}$ at $t_{\star}$ for different disorder strength $\Theta$. The upper bound in Sec. V. A. predicts a distribution around $\theta=\pi/2$. We see that $\theta_{i}$ is indeed distributed around $\pi/2$ for small disorder $\Theta\leq\pi/2$. However, the center of the distribution shifts to the left for larger $\Theta$. Both the finite-width distribution of $\theta_{i}$ and the finite-width distribution in the $\phi$ direction in Fig. 8 contribute to the discrepancy between the upper bound $\mathbb{R}_{<}$ and the numerical data. Simulations use $N=400$ atoms and $10^{3}$ QSDMF trajectories.
Figure S5: In complement to Fig. 11, (a) Error of DTWA for $N=5,10,15$ atoms coupled with one lossy cavity, (b) Burst time $t_{\star}$ vs. NN for atoms coupled with two cavities, for $\kappa/(\gamma N)=1,2,5,10,50$ (from bottom to top), and (c) The roughly linear relation between $\gamma t_{\star}N\ln(N)$ and $\gamma N/\kappa$ for $N=400$. $t_{\star}\sim\ln(N)/(\gamma N)$ scaling persists in the weakly non-Markovian regime, with the prefactor decreasing to 11 as $\kappa$ increases. Simulations use $10^{3}$ trajectories.
K Reduction of the SDE for Efficient Treatment of Cavity Dynamics
$\begin{split}d\alpha^{(0)}_{\text{R}/\text{L}}&=-\frac{\kappa}{2}\alpha^{(0)}_{\text{R}/\text{L}}dt+\sqrt{\frac{\kappa}{2}}dW_{\text{R}/\text{L}},\end{split}$ (S1)
and a deterministic part
[…]
For the initial state, we can let $\alpha^{(0)}_{\text{R}/\text{L}}(0)=\alpha_{\text{R}/\text{L}}(0)$ and $\alpha_{\text{R}/\text{L}}^{(1)}(0)=0$. In numerical simulations, the stochastic part (S1) can be solved using the Euler-Maruyama algorithm with a very small time step, to generate $\mathcal{N}$ stochastic trajectories of $\alpha^{(0)}_{\text{R}/\text{L}}$ to high precision. Then, for each trajectory of $\alpha^{(0)}_{\text{R}/\text{L}}$, we can input it as dynamical variables into the set of deterministic EOMs of spin variables in Eq. (9) and variables $\alpha^{(1)}_{\text{R}/\text{L}}$ in (S2). The resulting deterministic ordinary differential equations (ODE) can then be solved to high precision with higher-order integrators. This can make the higher-order integrators of ODE directly applicable to the dynamics of the spin variables.
L Irrelevance of Detuning in DTWA
In Sec. VI.1, we show the numerical results that detuning is irrelevant as $N$ gets sufficiently large in DTWA simulations. The argument there uses the assumption that the cumulant expansion of finite order is accurate. Here we show that, since each trajectory in DTWA is treated at the MF level, no such additional assumption is needed in this analysis. In the DTWA formalism, after the unitary transformation for the rotating frame, the EOMs (54) in Appendix B become
@@
with $\tilde{s}_{j}^{\pm}(t)=s_{j}^{\pm}(t)e^{\mp i\delta\omega_{j}t}$. As expected, this is just the Heisenberg EOM (VI.1) plus noises. Similar to Sec. VI.1, the detuning gives corrections $\sim\Delta_{\omega}t\sim(\Delta_{\omega}/\gamma)\ln(N)/N$ to the EOMs, which become negligible for $N/\ln(N)\gg\Delta_{\omega}/\gamma$. Since DTWA is accurate for this problem, this also indicates that the required cumulant order $C\ll N/\ln(N)$ for superradiance. The analysis using DTWA and that using CE mutually corroborate the conclusion that finite detuning is irrelevant for large $N$ in superradiance.